# Social Choice for AI Alignment: Dealing with Diverse Human Feedback

### v1

@@ 7 Which Traditional Social-choice-theoretic Concepts are Most Relevant?
A wide variety of concepts is studied in social choice. In general, the relevance of those concepts depends on the specific application. For example, consider the concept of false-name-proofness (Yokoo et al., 2001, 2004; Conitzer & Yokoo, 2010), which means that no participant can benefit from participating multiple times under multiple accounts. This concept is relevant when voting over the internet, but entirely irrelevant in, say, an in-person faculty meeting where faculty vote publicly by raising their hands.
@@ 7.1 Independence of Clones
Sometimes the introduction of a clone can affect the outcome of an election. For example, suppose a group of people are voting over where to go for dinner, and the only two alternatives are a Chinese restaurant and an Indian restaurant. 52% of the voters prefer the Chinese restaurant. But then, someone points out that the Chinese restaurant has two floors and argues that the two floors should be considered separate options. So now the alternatives are $C_{1}$, $C_{2}$, and $I$. It turns out nobody really cares all that much about the floor, but suppose that 26% of the voters prefer $C_{1}\succ C_{2}\succ I$, and 26% of the voters prefer $C_{2}\succ C_{1}\succ I$ (adding up to the original 52%). Further suppose that the voting rule used is Plurality, in which the alternative that appears at the very top of voters’ rankings the most often wins. [^5] Given that only the top-ranked alternative matters, in practical implementations of this rule, we simply ask voters to list only that alternative—the alternative the voter votes for. This results in the Indian restaurant now actually winning with 48% of the vote. This seems like an undesirable property for a voting rule to have; it would be better for the introduction of a clone never to make a difference. [^6] More precisely, introducing a clone should not affect whether a non-clone (e.g., the Indian restaurant in our example) is selected or which non-clone is selected. But it may affect which clone, if any, is selected (e.g., a clone-independent rule could select $C_{1}$ over $C_{2}$ in our example, if among the 48% of people who prefer $I$, a strict majority of them prefer $C_{1}$ to $C_{2}$). This latter desirable property is called independence of clones. Perhaps when choosing restaurants, this is not that important, as restaurants will rarely be clones (unless the floors of restaurants are treated separately). On the other hand, when choosing responses for a chatbot, it may be quite common for two responses to be very close to each other. This suggests that this property is more important in this context.
@@ 7.2 Strategic voting
What form might strategic voting in a context such as RLHF take? If rating responses on a scale from (say) 0 to 10, a natural strategy is to overreport. E.g., if one evaluator does not really like a response (at the level of a 3), but suspects that others would like it (say, two other evaluators that give a 6), then this evaluator may strategically give a rating of 0 to “compensate” for the other reviewers. This manipulation would be successful if we eventually aggregate ratings by taking their average: the average will be pulled down to 4, instead of the 5 that would result from reporting truthfully, so that the average is closer to the 3 that the evaluator believes is ideal. If instead we use the median as the aggregate, then this manipulation is ineffective—the median would remain 7. Indeed, the median is strategy-proof in this context: misreporting one’s preferences never helps, as long as one’s only goal is to move the median rating closer to one’s “true” rating.
@@ 7.3 Anonymity
In democratic contexts, a standard desideratum on voting rules is anonymity: if two voters swap their ballots before submitting them, the output of the voting rule will not change (the rules in Figure 1 all satisfy anonymity in this sense). This captures the idea that the voting rule should not favor some voters over others. Anonymity not only prohibits the extremes of dictatorship (recall Section 2.5), but even any kind of weighted voting wherein some voters’ votes count for more than others. However, in the context of AI development, one might consider aggregating human feedback in a way that violates anonymity (cf. the weighted majority rule discussed in Nitzan & Paroush 1982). Perhaps some evaluators are more experienced or more highly rated than others; perhaps some are influenced by others, so their input should not be considered completely independent inputs for aggregation; and so on. In general, whether the same democratic norms applied to voting also apply in an AI context is an important question for discussion.
@@ 7.4 Principles as Voters
While it is standard in social choice for the voters to be human agents, this is not the only interpretation of the mathematical framework of social choice theory. In some applications of social choice to AI ethics and safety, possibly including Constitutional AI (recall Section 2.4), we might regard different ethical principles as the “voters” who can rank or otherwise evaluate the outputs of an AI system. (cf. Greene et al. 2016.) This is analogous to applications of social choice theory in the philosophy of science, where the “voters” are theoretical virtues that may rank scientific theories differently (Okasha, 2011), or to multi-criteria decision-making, where the “voters” are relevant factors that may rank the options differently (Arrow & Raynaud, 1986). Of course, such ethical principles could themselves be outputs of some prior social choice procedure in which the voters are humans (cf. Collective Constitutional AI in Ganguli et al. 2023).
This principles as voters idea suggests a possible alternative architecture for applying social choice to AI—one sitting somewhere between the extremes of a spectrum that ranges from Constitutional AI at one end (in which principles are the whole show, while social choice does not appear) to the RLHF version of reinforcement learning as described above (in which principles play no role at all). In this alternative model, each respondent would be required to justify her rankings of alternative AI responses in terms of their level of satisfaction of each of a number of principles taken from a fixed menu. The AI system would use the results to train for several independent tasks: for each principle, separately learn how to rate responses to queries based on that principle alone; and learn how to aggregate those separate ratings into an overall rating of the responses. These would be composed to form the final stage of a simulated collective decision—the stage in which the voters are the principles.
@@ 8 How Should We Account for Behavioral Aspects and Human Cognitive Structures?
The theory of preference elicitation tends to be based on idealized assumptions; for example, each agent whose preferences are being elicited has well-defined and consistent preferences, and the agent answers in a way that is perfectly consistent with those preferences (or perhaps random noise is added). In reality, when we elicit human preferences, myriad behavioral effects kick in.
@@
This leads to a variety of questions about how to align AI systems in the context of such behavioral effects. Should we correct for the behavioral effect? That would seem to require having a model where such a behavioral effect obscures humans’ “true” values. But do these “true” values correspond to anything real in the world? Do we run the risk of the “correction” actually removing valuable information? Could the ability to make such “corrections” in fact be abused to intentionally remove inconvenient feedback?
9 How to Navigate a Multiplicity of AIs?
@@
Consider the example of a group of people voting over the restaurant where they will go for dinner. If there is significant disagreement in the votes, rather than forcing a minority to go to a restaurant that they really do not like, it can make sense to split the people into multiple groups, each going to their own restaurant. Similarly, in the current context, perhaps it makes sense to create multiple AI systems; for example, to recognize strong inter- and intra-cultural variations that have been identified in some non-homogenous populations (Awad et al., 2018; Peters & Carman, 2024). Depending on the situation, the people providing feedback might be split into groups ex ante (for example, country A makes one system based on the feedback of A’s citizens and country B another based on the feedback of B’s citizens), but also ex post, where we first collect feedback and then consider which people it makes sense to group together. The latter approach is closely related to the topic of representation in voting theory (Faliszewski et al., 2017).
@@
There is also the slightly different scenario where one AI system is in place, and some group of people believes that it is not serving them well. Hence, they might decide to pool their resources and create their own system. The literature on cooperative game theory (cf. Chalkiadakis et al. 2011), sometimes referred to as coalitional game theory, touches on these considerations (and indeed also plays a role in questions of representation, Aziz et al. 2017).
@@
Finally, let us highlight possible shortcomings to creating multiple AI systems. As in the restaurant example, it may have the result of unnecessarily dividing people into separate groups. Moreover, splitting into groups may not be feasible if it does not dovetail with existing social structures. For example, the US Federal Government may want to adopt a single system that will impact all its citizens, and adopting two systems would be tantamount to splitting the country in two. Finally, unlike in the case of the restaurants, the multiple AI systems may have to interact with each other, creating the risk of conflict between AIs with different goals. The nascent literature on cooperative AI (Dafoe et al., 2021; Conitzer & Oesterheld, 2023) may help keep these kinds of interactions from going horribly wrong. Nonetheless, it might be best to see if we can completely avoid having multiple AIs with competing goals, or at least design them in a way that makes conflict between them less likely.
@@ 10 Conclusion
Needless to say, the questions considered above are multifaceted and, as such, cannot be adequately addressed without complementary (not necessarily AI-specific) research. How best to make practical decisions, as well as associated legal and political considerations, provide further important avenues for future research.
@@ Acknowledgements
Impact Statement
As we briefly discussed in the introduction, there is perhaps a concern that feedback from a broader set of participants is more likely to be inconsistent and that consequently the resulting system will behave erratically. The idea that naïve aggregation of votes or judgments leads to inconsistency is a familiar one from social choice theory. For example, if three voters rank three alternatives $A$, $B$, and $C$ respectively as follows: $A\succ_{1}B\succ_{1}C$, $C\succ_{2}A\succ_{2}B$, and $B\succ_{3}C\succ_{3}A$, then a majority of voters prefers $A$ to $B$, a majority prefers $B$ to $C$, and a majority prefers $C$ to $A$. This illustrates that majority preferences are cyclical and thus arguably irrational. (See Figure 1 for another example.) We encounter similar issues in judgment aggregation (for an overview, see Endriss 2015). To illustrate this in our own context, say that it is broadly agreed that an output should be given if and only if it is both safe and helpful. Suppose evaluator 1 believes the output is safe but not helpful, and therefore should not be given. Evaluator 2 believes the output is helpful but not safe, and therefore should not be given. Evaluator 3 believes the output is both safe and helpful, and therefore should be given. Then a majority believes that the output is safe, a majority believes that it is helpful, but a majority believes that it should not be given—so that majority judgments are logically inconsistent. However, social choice theory is precisely concerned with how we should actually obtain consistent aggregations and therefore is well placed to address this issue. For example, one common strategy is to restrict to rational or consistent outputs only and among these find one that is in some sense “closest” to the reports (see, e.g., Elkind & Slinko 2015). Therefore, social choice theory is actually well positioned to help with the issue of inconsistencies from aggregation.

### v2

@@ 7 Which Traditional Social-Choice-Theoretic Concepts Are Most Relevant?
Social choice studies a wide variety of concepts, the relevance of which depends on the specific application. For example, consider the concept of false-name-proofness (Yokoo et al., 2001, 2004; Conitzer & Yokoo, 2010), meaning that no-one can benefit from participating multiple times under multiple accounts. This is relevant when voting over the internet, but irrelevant for in-person faculty meetings where faculty vote publicly by raising their hands.
@@ 7.1 Independence of Clones
Sometimes the introduction of a clone can affect the outcome of an election. Suppose a group of people are voting over where to go for dinner, and the only two alternatives are a Chinese restaurant and an Indian restaurant. 52% of the voters prefer the Chinese restaurant. But then, someone points out that the Chinese restaurant has two floors and argues that the two floors should be considered separate options. So now the alternatives are $C_{1}$, $C_{2}$, and $I$. Nobody really cares all that much about the floor, but suppose that 26% of the voters prefer $C_{1}\succ C_{2}\succ I$, and 26% of the voters prefer $C_{2}\succ C_{1}\succ I$ (adding up to the original 52%). Further suppose that the voting rule used is Plurality, in which the alternative that appears at the very top of voters’ rankings the most often wins. Then the Indian restaurant ends up winning now with 48% of the vote. This seems like an undesirable property for a voting rule to have; it would be better for the introduction of a clone never to make a difference. [^4] More precisely, introducing a clone should not affect whether a non-clone (e.g., the Indian restaurant in our example) is selected or which non-clone is selected. But it may affect which clone, if any, is selected. For instance, a clone-independent rule could select $C_{1}$ over $C_{2}$ in our example, if among the 48% of people who prefer $I$, a strict majority of them prefer $C_{1}$ to $C_{2}$. This latter desirable property is called independence of clones. Perhaps when choosing restaurants, this is not that relevant, as restaurants will rarely be clones (unless the floors of restaurants are treated separately). On the other hand, when choosing responses for a chatbot, it may be quite common for two responses to be very close to each other, indicating its importance to this context. We note that Borda Count (see Figure 1), which is implicitly used in some standard approaches to RLHF (Siththaranjan et al. 2023), badly violates independence of clones.
@@ 7.2 Strategic Voting
What form might strategic voting in a context such as RLHF take? If rating responses range on a scale from (say) 0 to 10, a natural strategy is to overreport. E.g., if one evaluator does not really like a response (at the level of a 3), but suspects that others would like it (say, two other evaluators give it a 6), then this evaluator may strategically give a rating of 0 to “compensate” for the other reviewers. This manipulation would be successful if we eventually aggregate ratings by taking their average: the average will be pulled down to 4, instead of the 5 that would result from reporting truthfully, so that the average is closer to the 3 that the evaluator believes is ideal. If instead we use the median as the aggregate, then this manipulation is ineffective—the median would remain 6. Indeed, the median is strategy-proof in this context: misreporting one’s preferences never helps, as long as one’s only goal is to move the median rating closer to one’s “true” rating (cf. Moulin 1980).
@@ 7.3 Anonymity
In democratic contexts, a standard desideratum on voting rules is anonymity: if two voters swap their ballots before submitting them, the output of the voting rule will not change (the rules in Figure 1 all satisfy anonymity in this sense). This captures the idea that the voting rule should not favor some voters over others. Anonymity prohibits not only the extremes of dictatorship (recall Section 2.5) but even any kind of weighted voting wherein some voters’ votes count for more than others. However, in the context of AI development, one might consider aggregating human feedback in a way that violates anonymity (cf. the weighted majority rule discussed in Nitzan & Paroush 1982). Perhaps some evaluators are more experienced or more highly rated than others; perhaps some are influenced by others, so their input should not be considered completely independent inputs for aggregation; and so on. In general, whether the same democratic norms applied to voting also apply in an AI context is an important question for discussion.
@@ 7.4 Principles as Voters
While it is standard in social choice for the voters to be humans, this is not required by the social choice theory framework. In some applications of social choice to AI ethics and safety, possibly including Constitutional AI (recall Section 2.4), we might regard different ethical principles as the “voters” who can rank or otherwise evaluate the outputs of an AI system (cf. Greene et al. 2016). This is analogous to applications of social choice theory in the philosophy of science, where the “voters” are theoretical virtues that may rank scientific theories differently (Okasha, 2011), or to multi-criteria decision-making, where the “voters” are relevant factors that may rank the options differently (Arrow & Raynaud, 1986). Of course, such ethical principles could themselves be outputs of some prior social choice procedure in which the voters are humans (cf. Collective Constitutional AI in Ganguli et al. 2023).
This suggests a possible alternative architecture for applying social choice to AI—one sitting somewhere between the extremes of a spectrum that ranges from Constitutional AI at one end (in which principles are the whole show, while social choice does not appear) to the RLHF version of reinforcement learning as described in previous sections (where principles play no role at all). In this alternative model, each respondent would be required to justify her rankings of alternative AI responses in terms of their level of satisfaction of each of a number of principles taken from a fixed menu. The AI system would use the results to train for several independent tasks: for each principle, separately learn how to rate responses to queries based on that principle alone; and learn how to aggregate those separate ratings into an overall rating of the responses. These would be composed to form the final stage of a simulated collective decision—the stage in which the voters are the principles.
@@ 8 How Should We Account for Behavioral Aspects and Human Cognitive Structures?
Preference elicitation often makes idealized assumptions, e.g., that each queried individual has well-defined and consistent preferences and will answer in a way that is consistent with them (up to some random noise). But in reality, myriad behavioral effects kick in. For example, McElfresh et al. (2021) study how (in the context of kidney allocation) humans can become indecisive when asked to give input that will have moral implications.
@@
This leads to a variety of questions about how to align AI systems in the context of behavioral effects. Should we correct for them? That would seem to require having a model where such a behavioral effect obscures humans’ “true” values. But do these “true” values correspond to anything real in the world? Do we run the risk of the “correction” actually removing valuable information? Could the ability to make such “corrections” in fact be abused to intentionally remove inconvenient feedback?
9 How Do We Navigate a Multiplicity of AIs?
@@
Consider, again, a group of people voting on a restaurant for their dinner. If there is a significant disagreement, then rather than forcing a minority to go someplace they really do not like, it can make sense to split into multiple groups, each going to their favorite restaurant. Similarly, perhaps it makes sense to create multiple AI systems—for example, to recognize strong inter- and intra-cultural variations that have been identified in some non-homogenous populations (Awad et al., 2018; Peters & Carman, 2024). Depending on the situation, the people providing feedback might be split into groups ex ante (e.g., each country makes a system based on their own citizens’ feedback), but also ex post, where we first collect feedback and then form groups of people. The latter approach is closely related to the topic of representation in voting theory (Faliszewski et al., 2017).
@@
There is also the slightly different scenario where one AI system is in place, and some group of people believe that it is not serving them well. Hence, they might decide to pool their resources and create their own system. The literature on cooperative game theory (cf. Chalkiadakis et al. 2011), sometimes referred to as coalitional game theory, touches on these considerations (and indeed also plays a role in questions of representation, as in Aziz et al. 2017).
@@
Finally, let us highlight possible shortcomings to creating multiple AI systems. As in the restaurant example, it may have the result of unnecessarily dividing people into separate groups. Moreover, splitting into groups may not be feasible if it does not dovetail with existing social structures. For example, the US Federal Government may want to adopt a single system that will impact all its citizens, and adopting two systems would be tantamount to splitting the country in two. Finally, unlike in the case of the restaurants, the multiple AI systems may have to interact with each other, creating the risk of conflict between AIs with different goals. The nascent literature on cooperative AI (Dafoe et al., 2021; Conitzer & Oesterheld, 2023) may help ensure these interactions do not result in adverse outcomes. Nonetheless, it might be best to see if we can completely avoid having multiple AIs with competing goals, or at least design them in a way that makes conflict between them less likely.
@@ 10 Conclusion
Needless to say, the questions considered are multifaceted and cannot be adequately addressed without complementary (and not necessarily AI-specific) research. The involved practical decisions and associated legal and political considerations are important topics for future research as well.
@@
Impact Statement
As we briefly discussed in the introduction, there is perhaps a concern that feedback from a broader set of participants is more likely to be inconsistent and that consequently the resulting system will behave erratically. The idea that naïve aggregation of votes or judgments leads to inconsistency is a familiar one from social choice theory. For example, if three voters rank three alternatives $A$, $B$, and $C$ respectively as follows: $A\succ_{1}B\succ_{1}C$, $C\succ_{2}A\succ_{2}B$, and $B\succ_{3}C\succ_{3}A$, then a majority of voters prefers $A$ to $B$, a majority prefers $B$ to $C$, and a majority prefers $C$ to $A$. This illustrates that majority preferences are cyclical and thus arguably irrational. (See Figure 1 for another example.) We encounter similar issues in judgment aggregation (for an overview, see Endriss 2015). To illustrate this in our own context, say that it is broadly agreed that an output should be given if and only if it is both safe and helpful. Suppose evaluator 1 believes the output is safe but not helpful, and therefore should not be given. Evaluator 2 believes the output is helpful but not safe, and therefore should not be given. Evaluator 3 believes the output is both safe and helpful, and therefore should be given. Then a majority believes that the output is safe, a majority believes that it is helpful, but a majority believes that it should not be given—so that majority judgments are logically inconsistent. However, social choice theory is precisely concerned with how we should actually obtain consistent aggregations and is therefore well placed to address this issue. For example, one common strategy is to restrict to rational or consistent outputs only and among these find one that is in some sense “closest” to the reports (see, e.g., Elkind & Slinko 2015). Therefore, social choice theory is actually well positioned to help with the issue of inconsistencies from aggregation.